OpenStudy (anonymous):

HELP SOLVING THE EQUATION: (-2 &1/7)p=-3

OpenStudy (anonymous):

Is your eq like this \[ -2 \frac{1}{7}p=-3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If yes then \[-2 \frac{1}{7}p=-3 \rightarrow \(\frac{-2 \times 7+1}{7})p=-3\] \[(\frac{-14+1}{7})p=-3 \rightarrow (\frac{-13}{7})p=-3 \rightarrow 13p=3 \times 7\] \[ 13p=21 \rightarrow p = \frac{21}{13}\]

OpenStudy (anonymous):

my answer choices have a whole number 1 or 6 so idk ?

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

no, my answer choices arent the answer you given me. The answer choices have either a whole number 1 or whole number 6

OpenStudy (anonymous):

\[-(2\frac{1}{7})p=-3\] \[-(\frac{15}{7})p=-3\] \[p=\frac{21}{15}=\frac{7}{5}\]

OpenStudy (anonymous):

\[1\frac{2}{5}\]